Two times a first number decreased by a second number is 4. The first number increased by twice the second number is 37. Find the numbers
Answer by scott8148(6628)

"Two times a first number decreased by a second number is 4" ___ 2f - s = 4

"The first number increased by twice the second number is 37" ___ f + 2s = 37

multiplying 1st eqn by 2 ___ 4f - 2s = 8

adding 2nd eqn (eliminating s)___ 5f = 45 ___ f = 9

substituting to find s ___ 2(9) - s = 4 ___ 14 = s